OpenStudy (anonymous):

without using graphs as an aid, use your knowledge of differentials to discuss the continuity of the function g(x)= 5x^3=10 at point x=4

OpenStudy (anonymous):

http://www.youtube.com/watch?v=dQw4w9WgXcQ this will help

OpenStudy (jack1):

5x^3=10 at point x=4 ... this doesn't work, sorry? 5x^3=10 therefore x = a constant... = cube root of 2... so x never equals 4

OpenStudy (anonymous):

Sorry I meant 5x^3-19 at point x=4

OpenStudy (jack1):

my first thing is always to "solve" for y: g(x) = 5x^3-19 g(4) = 5(4)^3-19 g(4) = 320-19 g(4) = 301 so at x = 4, y = 301 now, slope of the line (first derivative): g(x) = 5x^3-19 g'(x) = 15x^2 so at x = 4 g'(4) = 15(4)^2 g'(4) = 15 * 16 g'(4) = 240 therefore: slope is positive and crazy steep! now: is slope increasing or decreasing (second derivative): g(x) = 5x^3-19 g'(x) = 15x^2 g''(x) = 30x so at x = 4 g''(x) = 30x g''(4) = 30(4) g''(x) = 120 slope of the line is increasing at this point, and sharply increasing that is...

OpenStudy (anonymous):

ok, so my question then is for the actual answer, what is it? I have already figured out most of what you had written, I do not understand what to put down as an answer

OpenStudy (jack1):

 the question just says discuss the continuity of this function around the point 4, so not really sure, but would go with something like: function is positive in the cartesion plane at this point, function's slope is steep and positive at this point function continues to get steeper continuing past this point as x increases